OpenStudy (anonymous):

(√23 + √13i)(√23 – √13i)

OpenStudy (anonymous):

(a+b)(a-b)=a^2-b^2

OpenStudy (anonymous):

with complex numbers \[(a+bi)(a-bi)=a^2+b^2\]